### v2

# Integrating Knowledge Retrieval and Large Language Models for Clinical Report Correction

Abstract
This study proposes an approach for error correction in radiology reports, leveraging large language models (LLMs) and retrieval-augmented generation (RAG) techniques. The proposed framework employs a novel internal+external retrieval mechanism to extract relevant medical entities and relations from the report of interest and an external knowledge source. A three-stage inference process is introduced, decomposing the task into error detection, localization, and correction subtasks, which enhances the explainability and performance of the system. The effectiveness of the approach is evaluated using a benchmark dataset created by corrupting real-world radiology reports with realistic errors, guided by domain experts. Experimental results demonstrate the benefits of the proposed methods, with the combination of internal and external retrieval significantly improving the accuracy of error detection, localization, and correction across various state-of-the-art LLMs. The findings contribute to the development of more robust and reliable error correction systems for clinical documentation.
Integrating Knowledge Retrieval and Large Language Models for Clinical Report Correction
Jinge Wu1thanks: These authors contributed equally to this work., Zhaolong Wu211footnotemark: 1, Ruizhe Li3, Abul Hasan1, Yunsoo Kim1, Jason P.Y. Cheung2thanks: Corresponding authors., Teng Zhang [^2] footnotemark: 2, Honghan Wu1422footnotemark: 2 1University College London 2University of Hong Kong 3University of Aberdeen 4University of Glasgow {jinge.wu.20, honghan.wu}@ucl.ac.uk
1 Introduction
The rise of large language models (LLMs) have brought significant advancements to healthcare, with one of the fastest-growing applications being the generation of clinical notes Abacha et al. (2023a); Zhou et al. (2023); Rajpurkar et al. (2022). However, guaranteeing the quality and accuracy of these AI-generated notes remains a challenge, as LLMs may occasionally produce hallucinated content that deviates from factual information Ji et al. (2023); Guerreiro et al. (2023). Similarly, reports manually written by radiologists are not immune to errors, underscoring the need for meticulous error correction to achieve accurate clinical documentation, which is paramount for ensuring effective communication among healthcare professionals and ultimately, optimal patient care Wu et al. (2022); Brady (2017); Brady et al. (2021).
Current research on clinical report error correction is quite limited. The MEDIQA-CORR 2024 shared task Ben Abacha et al. (2024) attempts to evaluate the potential of using LLMs as solutions to locate and correct medical errors within clinical notesAchiam et al. (2023); Touvron et al. (2023). However, their study is limited by the use of synthetic data rather than real-world clinical notes, reducing its practicality in clinical settings. Furthermore, the task of error correction in clinical reports is inherently complex, requiring a large language model with strong command capabilities, as well as extensive clinical knowledge and reasoning skills.
In a clinical setting with limited computational resources, training a large, instruction-fine-tuned model specifically for this task may be impractical. To address this challenge, we propose a more feasible approach that guides the model to better perform the task by adjusting the instructions and enhancing context extraction. Our method leverages the strengths of existing LLMs while mitigating their limitations through a three-stage framework and retrieval-augmented generation (RAG) techniques.
The three-stage approach decomposes the error correction process into distinct subtasks: error detection, error localization, and error correction. By breaking down the complex task into more manageable and interpretable steps, we aim to improve the overall performance and explainability of the system, enabling a more comprehensive understanding of the error correction pipeline. This approach allows the model to focus on each subtask independently, leveraging the strengths of LLMs while mitigating their inherent limitations in the context of medical error detection and correction.
Furthermore, we introduce a retrieval-augmented generation (RAG) approach for clinical report error correction, which retrieves medical entities and relations both internally and externally to facilitate contextual and comprehensive reasoning. The purpose of internal retrieval is to detect inconsistencies within the provided report, whereas external retrieval seeks to augment the reasoning process when the given report lacks adequate contextual information.
In summary, our contributions are as follows:
  • We propose a novel framework that employs a retrieval-augmented generation approach for clinical report error correction. This framework retrieves medical entities and relations both internally and externally to facilitate contextual and comprehensive reasoning.
  • We introduce a three-stage approach for error detection and correction, which decomposes the task into distinct subtasks: error detection, error localization, and error correction. This decomposition not only serves as a benchmark but also enhances the explainability of the overall process.
  • We establish a benchmark for the clinical report error detection task by intentionally introducing common mistakes into real-world electronic health records, with the guidance of domain experts. This benchmark enables a rigorous evaluation of error detection capabilities.
  • We conduct thorough experiments to assess the effectiveness of our proposed approaches, demonstrating their performance through comprehensive evaluations and analyses.
2 Related Work
2.1 Retrieval-Augmented Generation
Retrieval-augmented generation (RAG) is a technique that enhances the performance of LLMs by integrating an external knowledge retrieval component Gao et al. (2023). The RAG architecture consists of two main modules: a retriever and a generator. The retriever is responsible for searching and ranking relevant information from a large-scale knowledge base, given an input query. The generator, usually a pre-trained LLM, takes the input query and the retrieved knowledge as context to generate the final output. RAG has shown promising results in various natural language processing tasks, such as question answering, fact verification, and dialogue systems Chen et al. (2024); Jin et al. (2024).
Traditional RAG systems often rely on unstructured text retrieval from large-scale corpora, such as Wikipedia or domain-specific databases Ke et al. (2024); Zhao et al. (2024). While this approach has shown success in various natural language processing tasks, it has limitations in retrieving valid information from massive free text compared to knowledge graph (KG) based retrieval Wu et al. (2023). KGs often contain high-quality, curated knowledge that has been manually or semi-automatically extracted from reliable sources Sanmartin (2024). This curated knowledge tends to be more accurate and consistent compared to the potentially noisy and contradictory information found in unstructured text corpora Jiang et al. (2024). The structured nature of KGs also enables easier integration of domain-specific knowledge, such as medical ontologies or scientific taxonomies, which can enhance the performance of RAG systems in specialized domains.
However, constructing and maintaining large-scale, high-quality KGs can be challenging and resource-intensive. To address this issue, this work presents a simple yet efficient method for constructing KGs directly from the reference data used in RAG systems. By leveraging the information already available in the reference corpus, this approach eliminates the need for extensive manual curation or external knowledge sources. Furthermore, this method circumvents the requirement for additional fine-tuning of LLMs, which can be prohibitively expensive, especially in low-resource environments such as the medical domain Thirunavukarasu et al. (2023); Sahoo et al. (2024).
2.2 RadGraph
RadGraph Jain et al. (2021) is a tool for extracting clinical entities and relations from radiology reports. It introduces a carefully designed information extraction schema that aims to capture most clinically relevant information within a radiology report while enabling consistent and efficient annotation. The key aspects of RadGraph’s entity and relation extraction approach are as follows: For entity extraction, RadGraph defines four entity types: Anatomy, Observation: Definitely Present, Observation: Uncertain, and Observation: Definitely Absent. Anatomy refers to anatomical body parts mentioned in the report, while observations encompass words associated with visual features, physiological processes, or diagnostic classifications. This novel entity typology facilitates comprehensive extraction of clinically salient information. Regarding relation extraction, RadGraph employs three relation types: suggestive_of, located_at, and modify. Suggestive_of captures inferred relationships between two observations. Located_at links an observation to an anatomy, indicating the anatomical location or context of the observation. Modify encodes modifier relationships, whereby one entity quantifies or scopes another entity of the same type. This principled relation schema enables RadGraph to represent complex semantic relationships within radiology reports. Our work mainly builds on RadGraph to extract entities and relations from radiology reports.
3 Method
Figure 1 illustrates the proposed work, which integrates internal and external retrieval, rephrasal augmentation, and finally with a three-stage inference process for step-by-step inference. The internal and external retrieval extracts relevant information from the given note and reference corpus. Then the retrieved entities and relations are rephrased into sentences. The three-stage inference process then utilizes the retrieved information to detect, localize, and correct errors in the clinical note. By combining these components, the proposed system aims to improve the accuracy and quality of clinical reports.
Figure 1: An overview of our work. ANAT-DP: Anatomy-Definitely Present. OBS-DP: Observation-Definitely Present. OBS-DA: Observation-Definitely Absent.
3.1 Three-Stage Inference
To enhance the inference capabilities and explainability of error correction in clinical reports, we propose a three-stage approach inspired by Ben Abacha et al. (2024). This method decomposes the task into three distinct subtasks, allowing the model to focus on smaller, more manageable objectives, leading to improved performance and increased interpretability.
Our three-stage approach consists of:
  • Error Detection: The model identifies whether an error is present in the given clinical text. This binary classification task enables the model to distinguish between correct and incorrect information by learning the characteristics that indicate the presence of errors.
  • Error Localization: If an error is detected, the model pinpoints its precise location within the text. This stage involves a fine-grained analysis of the clinical report to identify the specific words, phrases, or sentences that contain the error.
  • Error Correction: Based on the information gathered from the previous stages, the model generates the corrected sentence. This stage leverages the knowledge of the error’s presence and location to create a contextually appropriate correction that maintains the overall coherence and accuracy of the clinical report.
This approach offers several advantages over end-to-end error correction. By decomposing the task, the model can learn more efficiently and effectively, capturing the intricacies of error detection and correction more accurately. It also enhances explainability by providing clear insights into the model’s thought process at each stage. Furthermore, this step-by-step method helps reduce hallucination and encourages more targeted, accurate corrections.
3.2 Internal and External Retrieval
Here we introduce a novel graph and knowledge-based retrieval method for prompt engineering to enhance our system’s reasoning capabilities across all phases. This approach simplifies complex information and eliminates redundancies, facilitating more efficient and effective reasoning. Our study focuses on error correction in radiology reports, specifically addressing the Findings and Impression sections. The Findings section documents radiographic observations, while the Impression section provides diagnostic conclusions based on these findings. Our retrieval technique incorporates both internal and external retrieval processes to enhance the accuracy and consistency of radiology report error correction. The internal retrieval transforms the report into a graph representation using RadGraph, focusing on key medical entities and their relationships, while the external retrieval leverages a large-scale dataset of clinical reports to provide supplementary knowledge and context for more robust error detection.
Internal Retrieval.
Inspired by Yan et al. (2023), we transform the report into a graph representation using RadGraph. In the resulting graph, nodes correspond to two main types of entities: anatomical entities (e.g., lungs, ribs) and observational entities (e.g., clear, acute). Edges connecting these nodes denote relations such as modification (an observational entity modifying an anatomical entity), location (specifying the anatomical location of an observation), or suggestion (an observation suggesting a particular diagnosis).
The extraction of these key medical mentions and their relationships is expected to enable more targeted reasoning by the model, as it can focus on the most salient information while considering the connections between various entities. Moreover, RadGraph provides additional contextual information by categorizing observations based on their presence or absence in the report. Observations are denoted as either “Observation: Definitely Present" (OBS-DP) or “Observation: Definitely Absent" (OBS-DA). This explicit labeling of the presence or absence of observations enhances the input for the model, allowing it to better identify contextual inconsistencies within the report. By leveraging this rich graph-based representation of the clinical report, our method aims to improve the accuracy and consistency of the model’s reasoning process.
External Retrieval.
In addition to retrieving information from the given report, our method employs an external retrieval process to extract relevant entities and knowledge from a large-scale dataset of clinical reports, which we refer to as reference data. The purpose of this external retrieval is to provide supplementary knowledge and clinical findings from other reports that can aid in identifying errors or inconsistencies in the current report. By leveraging a broader context of medical information, our method aims to improve the accuracy and robustness of the error detection process.
To ensure consistency in data representation, we apply RadGraph to extract entities and relations from the reference data in the same manner as we do for the given report. The external retrieval process allows our method to access a vast repository of medical knowledge and real-world clinical findings. This can be particularly beneficial in cases where the given report may contain rare or complex medical conditions that require additional context for accurate interpretation. By drawing upon the collective knowledge present in the reference data, our method can make more informed decisions and potentially identify errors that might be difficult to detect based solely on the information provided in the given report.
3.3 Rephrasal Augmentation
When applying RadGraph to extract entities and relations from radiology reports, we further rephrase the extracted information into concise sentences to enhance interpretability and downstream reasoning capabilities. RadGraph proposes a principled relation extraction schema to reconstruct radiology report findings from extracted clinical entities and their semantic relations. The schema defines three relation types: suggestive_of, located_at, and modify. These can be used to rephrase the entities and relations into sentences. For instance, consider a scenario where we have two relation triples: $\langle lower,modify,lobe\rangle$ and $\langle opacity,located\_at,lobe\rangle$. The first triple indicates that the entity lower modifies the entity lobe, while the second triple signifies that the entity opacity is located at the entity lobe. By combining these two triples based on their common entity lobe, we can generate a meaningful phrase: lower lobe opacity. This phrase captures the essential information conveyed by the individual relation triples, presenting it in a more concise and understandable manner. Furthermore, our method takes into account the negation of mentions, which is crucial for accurately representing the absence of certain findings. In cases where a mention is negated, we incorporate the word no at the beginning of the generated phrase. For example, if the relation triple $\langle opacity,located\_at,lobe\rangle$ is negated, we would generate the phrase no lobe opacity.
For a given report, our method retrieves the top $k$ most similar reports from the reference data to provide additional context and support the error detection process. These retrieved reports serve as extra information in the input, augmenting the knowledge available to the model.
4 Experiment
4.1 Dataset
Our study utilizes real-world clinical reports from the MIMIC-CXR dataset Johnson et al. (2019). MIMIC-CXR is a large-scale collection of ICU hospital data, providing a valuable resource for our analysis. From this dataset, we employ 112,251 radiology reports as reference data, representing correctly annotated ground truth reports. Additionally, we curate 1,622 radiology reports as evaluation data, where a subset of these reports are purposefully corrupted to simulate real-world annotation errors (see details in Table 1).
Reference data will be used for external retrieval to provide augmented knowledge and clinical findings for error detection and correction. Evaluation data, on the other hand, refers to the data used for evaluation of the task, where some of the reports contain errors. More data statistics can be found in Appendix (Table 10)
To introduce realistic errors into the evaluation data, we employ two strategies with the guidance of clinical experts and radiologists. Firstly, we replace essential clinical observations with irrelevant ones. Specifically, we utilize the 12 important observations from the CheXpert dataset Irvin et al. (2019) (excluding “No Finding" and “Support Devices" from the original 14 observations) and substitute them with unrelated observations carefully selected to be clinically irrelevant to the report context (see the lists in Table 8 and 9 in Appendix). Secondly, we selectively alter the polarity of observations by removing negation words such as “no" or “no evidence of", effectively transforming negated findings into affirmed ones. This process aims to mimic common annotation mistakes where the presence or absence of clinical findings is incorrectly recorded. By corrupting a subset of the evaluation data through these two strategies, we aim to create a more challenging and realistic benchmark that captures the types of errors that may arise in practical radiology report annotation scenarios.
Correct Incorrect Total
FIND IMPN
Reference Data 112,251 - - 112,251
Evaluation Data 512 582 528 1,622
Table 1: Data used in this work. FIND: Findings (# of reports that have errors in the Findings). IMPN: Impression (# of reports that have errors in the Impression). Each report can only have at most one error (either it is from Findings or Impression).
4.2 Evaluation
We report performance on the three-stage inference process, which includes error detection, localization, and correction. We calculate the accuracy for error detection and localizaiton. To evaluate the performance of the generative error correction stage, we use an aggregate Natural Language Generation (NLG) score, namely AggNLG, that combines ROUGE-1 Lin (2004), BERTScore Zhang et al. (2019), and BLEURT Sellam et al. (2020). This aggregate score has been shown to align best with human judgement compared to other NLG metrics, providing a comprehensive assessment of the quality and semantic similarity between the generated corrections and the ground truth Abacha et al. (2023b).
The calculation of error correction requires the model to analyze the input radiological report: if an error is detected, the model will correct it by generating the amended report entirety; otherwise, it will return “NA". The scoring process is conducted exclusively when both the model and the ground truth identify errors (neither being “NA"), and the rectified results are compared. Radiological report errors typically involve minor corrections of a few words, leaving the majority of the report content unaltered, which explains the generally higher NLG scores in the result tables.
4.3 Baseline
For comparison, we select various state-of-the-art LLMs from both the general and medical domains to benchmark our proposed approach. In the general domain, we choose the latest LLMs, which are LLaMA3 and Phi3, as our base modelsAbdin et al. (2024). These models have demonstrated strong performance across a wide range of natural language processing tasks. For LLaMA3, we utilize the 8B parameter version (LLaMA3-8B [^1] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct) due to computational resource limitations, opting out of the larger 70B model. For Phi3, we compare different model sizes, including Phi3-mini-3.8B [^2] https://huggingface.co/microsoft/Phi-3-mini-4k-instruct, Phi3-small-7B [^3] https://huggingface.co/microsoft/Phi-3-small-8k-instruct, and Phi3-medium-14B [^4] https://huggingface.co/microsoft/Phi-3-medium-4k-instruct, to investigate the impact of model scale on the performance. Furthermore, to explore the potential benefits of domain-specific knowledge, we consider two medical LLaMA3-8B variants, Llama3-Aloe-8B-Alpha [^5] https://huggingface.co/HPAI-BSC/Llama3-Aloe-8B-Alpha and MMedLM2666https://huggingface.co/Henrychur/MMedLM2Qiu et al. (2024); Gururajan et al. (2024). These models built on LLaMA3-8B, and have been further fine-tuned on a wide range of medical instructional datasets, synthetic medical data, medical textbooks, medical websites, etc.
5 Result
5.1 Comparisons with Three-Stage Inference
Table 2 presents a comparison of the performance of various language models on the tasks of error correction using three-stage inference and end-to-end approaches. The results demonstrate a clear trend: the three-stage inference method generally yields superior AggNLG scores compared to end-to-end approaches across most models examined.
MMedLM2 and Aloe, both medical language models, show the most substantial gains of 10.47% and 26.84% respectively. Interestingly, Phi3-mini exhibits a slight performance decrease (-0.28%), suggesting that the efficacy of this approach may depend on model characteristics. Larger size models, including Phi3 variants (small and medium) and Llama3-8B, demonstrate moderate to significant improvements ranging from 4.53% to 9.95%. This trend indicates a possible positive correlation between model scale and the effectiveness of the three-stage approach, potentially due to larger models’ enhanced ability to leverage task decomposition.
Overall, these findings suggest that decomposing the error correction process into distinct stages of detection, localization, and correction can lead to enhanced performance compared to end-to-end approaches. This improvement is likely attributable to the structured nature of the three-stage method, which allows models to focus on specific aspects of the task sequentially, potentially mitigating error propagation and enabling more precise corrections.
End-to-End Three-stage Inference
MMedLM2 47.80 58.27 (+10.47)
Aloe 63.33 90.17 (+26.84)
Phi3-mini 74.36 74.08 (-0.28)
Phi3-small 80.03 86.57 (+6.53)
Phi3-medium 84.47 90.25 (+5.79)
Llama3-8B 84.34 94.29 (+9.95)
Table 2: Comparisons with three-stage inference and end-to-end inference of error correction. AggNLG is the averagte score of ROUGE-1, BERTScore, and BLEURT.
5.2 Comparisons with Internal and External Retrieval
Table 3-5 present comparisons of internal and external retrieval methods on error detection (Table 3), error localization (Table 4), and error correction (Table 5).
Baseline Internal Internal + Ex
MMedLM2 41.49 39.33 73.05 (+31.56)
Aloe 45.31 35.14 67.26 (+21.95)
Phi3-mini 67.26 69.66 73.06 (+5.80)
Phi3-small 79.03 78.48 80.21 (+1.18)
Phi3-medium 73.67 68.93 79.04 (+5.37)
Llama3-8B 37.79 37.92 62.27 (+24.48)
Table 3: Comparisons of internal and external retrieval on error detection (acc%).
Table 3 shows the accuracy on error detection. All models demonstrate significant improvements when employing both internal and external retrieval. MMedLM2 exhibits the most substantial enhancement, with a 31.56% increase, while Phi3-small shows the least improvement at 1.18%. Notably, internal retrieval alone yields limited benefits, with only Phi3-mini and Llama3-8B showing marginal improvements.
Baseline Internal Internal + Ex
MMedLM2 30.94 29.28 46.05 (+15.11)
Aloe 43.34 34.83 51.35 (+8.01)
Phi3-mini 47.71 39.95 52.65 (+4.94)
Phi3-small 63.44 65.22 65.04 (+1.60)
Phi3-medium 69.73 65.65 63.44 (-6.29)
Llama3-8B 37.29 36.99 53.14 (+15.85)
Table 4: Comparisons of internal and external retrieval on error localization (acc%).
Table 4 illustrates error localization accuracy. All models show improvement while using both internal and external retrieval, except for Phi3-medium. Llama3-8B and MMedLM2 improve the most by more than 15%. Similar to error detection, it is noticed that internal retrieval does not provide effectiveness for this task as well, where only Phi3-small shows increased accuracy with internal retrieval.
Baseline Internal Internal + Ex
MMedLM2 58.27 52.66 53.50 (-4.77)
Aloe 90.17 89.81 74.77 (-15.40)
Phi3-mini 74.08 76.78 78.85 (+4.77)
Phi3-small 86.57 89.09 86.67 (+0.10)
Phi3-medium 90.25 92.41 92.25 (+2.00)
Llama3-8B 94.29 94.49 94.43 (+0.14)
Table 5: Comparisons of internal and external retrieval on error correction (AggNLG).
Table 5 shows the error correction results. The combination of internal and external retrieval further enhances the error correction performance for most models, though the two medical LLMs do not benefit from the retrieval. Phi3-mini achieves the highest relative improvement by 4.77%. Similarly, Phi3-small and Phi3-medium and Llama3-8B benefit from the addition of external retrieval, with their AggNLG scores rising to 86.67% and 92.25%, and 94.43% respectively. These results suggest that the effectiveness of external retrieval in enhancing error correction performance varies across different language models. General domain models, such as Phi3 and Llama3 benefit significantly from the combination of internal and external retrieval, however, medical LLMs (MMedLM2 and Aloe), may not experience substantial gains or even exhibit a slight performance decline. This indicates medical LLMs may need different retrieval strategies.
Error Detection (acc %) Error Localization (acc %) Error Correction (AggNLG)
Simple RAG Ours Simple RAG Ours Simple RAG Ours
MMedLM2 52.65 73.05 (+20.4) 35.88 46.05 (+10.17) 48.13 53.50 (+5.37)
Aloe 58.81 67.26 (+8.45) 44.51 51.35 (+6.84) 74.28 74.77 (+0.49)
Phi3-mini 65.59 73.06 (+7.47) 42.97 52.65 (+9.68) 75.76 78.85 (+3.09)
Phi3-small 72.51 80.21 (+7.7) 58.93 65.04 (+6.11) 84.56 86.67 (+2.11)
Phi3-medium 73.13 79.04 (+5.91) 60.91 63.44 (+2.53) 88.87 92.25 (+3.38)
Llama3-8B 51.84 62.27 (+10.43) 46.92 53.14 (+6.22) 88.34 94.43 (+6.09)
Table 6: Comparisons of our method with simple RAG. Our methods refer to the internal+external retrieval.
Simple RAG Ours
MMedLM2 30.51 22.0 -8.51
Aloe 25.6 16.1 -9.55
Phi3-mini 24.1 15.5 -8.58
Phi3-small 25.0 15.9 -9.11
Phi3-medium 31.4 21.9 -9.45
LLAMA3 8B 23.2 14.5 -8.68
Table 7: Inference efficiency comparison with simple RAG and our method (seconds taken per instance).
5.3 Comparison with Simple RAG
To further evaluate the effectiveness of our proposed method, we conducted a comparative analysis against a simple Retrieval-Augmented Generation (RAG) approach. While simple RAG typically involves retrieving relevant documents from a reference corpus and providing them as additional context to the language model, our method combines both internal and external retrieval mechanisms, focusing on key medical entities and their relationships.
Table 6 presents a comparison of our method with simple RAG across the tasks of error detection, localization, and correction. The results demonstrate that our method consistently outperforms simple RAG across all evaluation metrics. In error detection, our approach shows an average improvement of 10.06 percentage points over simple RAG. Notably, the MMedLM2 model exhibits the most significant enhancement, with an increase of 20.4 percentage points. For the error localization task, our method achieves an average improvement of 6.92 percentage points, with the Phi3-mini model showing the most substantial gain of 9.68 percentage points. In terms of error correction, our method consistently yields improvements in AggNLG scores. Although the magnitude of improvement is relatively smaller compared to the detection and localization tasks, all models demonstrate positive changes. The Llama3-8B model shows the most significant improvement, with an increase of 6.09 points in AggNLG score.
Table 7 compares the inference efficiency of our method against simple RAG. The results indicate that our method significantly outperforms simple RAG in terms of processing time per instance, with an average reduction of 9.15 seconds. This efficiency gain can be attributed to the more targeted information retrieval strategy in our method, enabling the model to locate and process relevant information more expeditiously.
These results underscore the advantages of our proposed retrieval method. By integrating internal and external retrieval mechanisms, our approach not only enhances the accuracy of error detection, localization, and correction but also significantly improves inference efficiency. This demonstrates that a well-designed retrieval strategy can lead to substantial performance improvements in the task of clinical report error correction.
The superior performance of our method can be attributed to several factors. Firstly, the graph-based internal retrieval allows for a more nuanced understanding of the report’s content, capturing intricate relationships between medical entities that might be overlooked in a simple text-based retrieval. Secondly, the external retrieval provides a broader medical context, potentially helping the model to identify subtle inconsistencies or errors that may not be apparent from the report alone.
6 Discussion
The results presented in this study demonstrate the effectiveness of the internal and external retrieval method in enhancing the performance of language models on error detection, localization, and correction tasks. Across all three tasks, our proposed approach consistently improves upon the baseline and internal retrieval methods for most of the models tested.
The two medical-specific models, MMedLM2 and Aloe, benefits the most with error detection but the least with error correction. It is noticeable that MMedLM2 performs the worst among all models. When provided with relevant information through the retrieval approach, Aloe demonstrates improved performance, while MMedLM2 continues to generate excessive and repetitive content. This suggests that fine-tuning may have impacted the language capabilities of the base Llama3 model. Overall, the fine-tuned models do not perform as well as the general-purpose Llama3 model.
One notable observation while looking into detail with the models’ predictions is the varying performance of the models based on their underlying architectures and training data. The Llama3-8B model exhibits a strong ability to follow instructions, performing well when provided with clear and detailed prompts. However, its performance may suffer when given ambiguous instructions. In contrast, the Phi3 models demonstrate extensive knowledge and perform well even with vague instructions. However, when the instructions conflict with their internal understanding or knowledge, their performance tends to decline. Additionally, the Phi3 models are less adept at following instructions compared to Llama3-8B, occasionally deviating from the required output format or generating excessive content.
Moreover, we also conduct human evaluation for further analysis. More detail can be found in the Appendix. Our human evaluation results (Figure 3) reveal that the three-stage inference (internal + external) method significantly outperforms other approaches across accuracy, factual consistency, and clinical relevance. This improvement is particularly notable for models like Phi3-medium and Llama3-8B, with MMedLM2 showing dramatic increases in factual consistency and clinical relevance. These findings underscore the importance of human assessment in evaluating medical report generation models and demonstrate the enhanced capabilities of our proposed method across various model architectures.
In conclusion, our three-stage inference process, integrated with external knowledge retrieval, proves to be a robust strategy for improving accuracy, factual consistency, and clinical relevance in medical text analysis. This approach significantly improves models’ reasoning capabilities and task-specific performance without domain-specific fine-tuning, thereby enhancing efficiency and explainability while preserving general language capabilities.
Future research should focus on refining retrieval mechanisms, optimizing external knowledge integration, and enhancing models’ utilization of summary information. These advancements will be crucial in further improving the reliability and versatility of language models in healthcare applications.
Limitation
While our current study focuses on radiology reports, we recognize the potential for extending our method to a broader range of clinical notes. In future work, we plan to expand the scope of our research by exploring the application of RadGraph to a wider variety of medical entities, going beyond radiographic studies. By incorporating entities and relations from diverse types of clinical notes, such as progress notes, discharge summaries, and consultation reports, we aim to generalize our error correction method to a more comprehensive set of medical documents.
The study also highlights the varying performance of LLMs based on their underlying architectures, training data, and instruction clarity. While general-purpose LLMs like Llama3-8B and Phi3 benefit from the proposed approach, domain-specific models, such as MMedLM2 and Aloe, exhibit limitations that warrant further investigation. It is suggested different retrieval strategies should be tailored for medical LLMs.
Furthermore, to obtain a more robust assessment of the error correction performance, enhanced evaluation methods are also should be considered in the future. One promising avenue is to incorporate human evaluations from domain experts, such as radiologists and other clinicians. These expert assessments would provide invaluable insights into the clinical relevance and accuracy of the generated corrections, complementing the automated evaluation metrics used in this study.
Ethics Statement
This work uses the de-identified clinical notes in MIMIC-CXR. We complete the Collaborative Institutional Training Initiative (CITI) Program’s “Data or Specimens Only Research” course [^7] https://physionet.org/content/shareclefehealth2013/view-required-trainings/1.0/#1 and sign the data use agreement to get access to the data. We strictly follow the guidelines and only use locally hosted LLMs with the data.
Appendix A Appendix
Experiment Setups.
All the experiments are conducted with 4 NVIDIA GeForce RTX 3090 24576MiB. For our LLM experiments, we utilize the AutoModelForCausalLM from the Hugging Face Transformers library. The model is loaded from the specified model_id, with the torch_dtype parameter set to torch.bfloat16. We set the max_new_tokens=300, do_sample=True, temperature=0.001 and top_p=0.8, all other hyperparameters remains unchanged as default values.
RAG Pipeline.
We use the e5-large-unsupervised model Wang et al. (2022) to transform internal and external datasets into a vectorized database. This process involves applying cosine similarity to find the most similar texts based on the input radiology reports. The experiment uses parameters set to chunk_size=1000 and chunk_overlap=100.
Evaluation Data Preparation.
Table 8 and 9 shows the words used for extracting and replacing in the original reports to construct the benchmark data. We got the list from radiologists where we chose 12 important observation entities to replace, and 41 words that is not relevant to the report as the replacement words.
Words to Replace
Atelectasis
Cardiomegaly
Consolidation
Edema
Enlarged Cardiomediastinum
Fracture
Lung Lesion
Lung Opacity
Pleural Effusion
Pleural Other
Pneumonia
Pneumothorax
Table 8: Words to replace
Replacement Words
Asthma; Costochondritis; Pulmonary Embolism;
Thoracic Outlet Syndrome; Tracheitis;
Tracheomalacia; Vocal Cord Dysfunction;
Pharyngitis; Laryngitis; Mesothelioma;
Obstructive Sleep Apnea; Aspergillosis;
Appendicitis; Gastroesophageal Reflux Disease;
Crohn’s Disease; Ulcerative Colitis; Gallstones;
Pancreatitis; Hepatitis; Cirrhosis; Peptic Ulcer;
Irritable Bowel Syndrome; Celiac Disease
Diverticulitis; Hemorrhoids; Anal Fissure
Intestinal Obstruction; Gastroparesis
Cholecystitis; Gastric Ulcer; Duodenal Ulcer
Esophageal Varices; Achalasia; Barrett’s Esophagus
Esophageal Cancer; Pancreatic Cancer;
Inflammatory Bowel Disease; Colorectal Cancer
Liver Cancer; Gastric Cancer; Hiatal Hernia;
Esophageal Stricture
Table 9: Replacement words.
Reference Data Evaluation Data
Findings (%) Impression(%) Findings(%) Impression(%)
ANAT-DP 1,261,293 (46) 306,632 (37) 21,090 (46) 5,475 (42)
OBS-DP 1,002,276 (37) 319,179 (39) 18,726 (41) 6,215 (48)
OBS-U 81,331 (3) 65,491 (8) 1,559 (3) 1,231 (9)
OBS-DA 371,429 (14) 135,573 (16) 4,859 (11) 136 (1)
Total Entities 2,716,329 (100) 826,875 (100) 46,234 (100) 13,057 (100)
Modify 1,124,901 (59) 333,435 (60) 20,568 (62) 5,754 (61)
Located_at 712,778 (38) 182,083 (33) 11,373 (34) 2,913 (31)
Suggestive_of 62,372 (3) 39,120 (7) 1,213 (4) 700 (7)
Total Relations 1,900,051 (100) 554,638 (100) 33,154 (100) 9,367 (100)
Table 10: Data Statistics. ANAT-DP: Anatomy-Definitely Present. OBS-DP: Observation-Definitely Present. OBS-U: Observation-Uncertain. OBS-DA: Observation-Definitely Absent.
Human Evaluation
Applying LLMs in the medical domain may not be entirely convincing if relying solely on automated evaluation metrics. This is because automatic evaluation metrics primarily focus on assessing semantic similarity and sentence structure, while in the medical field, even minor errors in generated text can have severe consequences. To provide a more comprehensive evaluation of the effectiveness of our methods, we conducted a human assessment one the generated tasks, which is the error correction task. We choose 50 examples with six models across the three different inference approaches for comparison: end-to-end, three-stage inference (baseline), and three-stage inference (Internal + Ex). This allowed us to evaluate the performance of the models in a more nuanced manner beyond relying solely on automated evaluation metrics. Detailed description of the evaluation criteria can be found in Table 11. We use a Likert scale ranging from 1 to 5 for the evaluation
Metrics
Accuracy: The model correctly identifies and addresses errors in the medical report.
- Does the model accurately identify all errors in the report?
- Are the corrections made by the model accurate and appropriate?
Factual Consistency: The model’s output maintains consistency with established medical facts and does not introduce incorrect information.
- Does the corrected information align with current medical knowledge?
- Are there any instances of contradictory or inconsistent information in the corrections?
Clinical Relevance: The model’s corrections and additions are clinically appropriate and enhance the overall quality and usefulness of the medical report.
- Do the corrections improve the clinical value of the report?
- Are the changes relevant to the patient’s condition and potential treatment?
Table 11: Metrics used for human evaluation.
Figure 3 presents the human evaluation scores for six models across three critical dimensions: Accuracy, Factual Consistency, and Clinical Relevance. The results reveal notable performance differences among the evaluated methods.
The three-stage inference (baseline) method consistently outperforms the end-to-end approach, albeit by a small margin. This slight edge is evident across all three dimensions for most models. During the human evaluation process, we observed that outputs from models using the three-stage inference (baseline) method exhibited more complete and coherent report structures. This observation may explain why this method’s automatic evaluation metrics were higher than those of the end-to-end approach, despite the relatively small difference in human evaluation scores.
Most strikingly, the three-stage inference (internal + external) method demonstrates significant improvements in model performance across all three dimensions. This method consistently outperforms other inference approaches, with all models scoring notably higher in accuracy, factual consistency, and clinical relevance. The improvement is particularly pronounced for models like Phi3-medium and Llama3-8B, where the gains are substantial across all metrics. Interestingly, the magnitude of improvement varies across models. For instance, the MMedLM2 model shows a dramatic increase in performance when using the three-stage inference (internal + external) method, especially in terms of factual consistency and clinical relevance. In contrast, the Aloe model exhibits a more modest, though still significant, improvement.
The reason why the automatic evaluation results did not show a highly significant improvement for the three-stage inference (internal + external) method compared to the three-stage inference (baseline) method may be that automatic evaluation metrics primarily focus on structure and overall semantic similarity. These metrics may not effectively capture the correct identification and correction of specific errors in the reports, which are crucial aspects that human evaluators can discern.
This discrepancy between automatic and human evaluation results underscores the importance of human assessment in evaluating the quality of medical report generation models. It suggests that while automatic metrics are useful for initial screening and large-scale evaluations, they may not fully capture the nuanced improvements in accuracy, factual consistency, and clinical relevance that are critical in medical contexts.
Case Study
Figure 2: Case study. Our method refers to the three-stage inference with internal + external retrieval. We highlight the correct part in red and the error part in blue.
The case study (Figure 2) demonstrates significant improvements in error detection and correction capabilities using our method across various models. Among the six models examined, Llama3-8b, Phi3-small, and Phi3-medium successfully corrected the error, showcasing the effectiveness of our approach. Notably, Phi3-medium’s correction perfectly matched the ground truth, while Llama3-8b and Phi3-small provided semantically correct corrections with slight variations in expression. Compared to the end-to-end approach, our method enabled Aloe and Phi3-mini to successfully identify the erroneous sentences, although they fell short in providing accurate corrections. This improvement in error detection, even without perfect correction, highlights the enhanced capabilities our method brings to these models.
Figure 3: Human evaluation comparison.
Figure 4: Prompt used for end-to-end and three-stage inference.
Figure 5: Prompt used for internal and external retrieval.